# Foundations of Reasoning in Language Models (FoRLM)

### v2

@@ Abstract
Test-time reasoning architectures such as those following the Generate–Verify paradigm – where a model iteratively refines or verifies its own generated outputs – prioritise generation and verification, but exclude the monitoring processes that determine when and how reasoning should begin. This omission may contribute to the prefix dominance trap, in which models commit early to suboptimal reasoning paths and seldom recover, yielding roughly 20% accuracy loss. We address this architectural gap by formalising Flavell’s and Nelson and Narens’ metacognitive theories into computational specifications, proposing the Monitor–Generate–Verify (MGV) framework. MGV extends the Generate–Verify paradigm by adding explicit monitoring that captures metacognitive experiences (from difficulty assessments to confidence judgements) before generation begins and refines future monitoring through verification feedback. Though we present no empirical validation, this work provides the first systematic computational translation of foundational metacognitive theories, offering a principled vocabulary for understanding reasoning system failures and suggesting specific architectural interventions for future test-time reasoning designs.
@@ 1 Introduction
Once language models commit to an initial reasoning strategy, subsequent verification rarely helps; this prefix dominance trap causes nearly 20% performance degradation when models choose suboptimal approaches, with virtually no recovery possible through refinement (Luo et al., 2025). Today’s dominant Generate-Verify (G-V) test-time reasoning architectures (Weng et al., 2023; Madaan et al., 2023; Lee et al., 2025; Zhang et al., 2024) exemplify this limitation through their very design. They operate through immediate generation followed by iterative refinement, without assessing task characteristics or selecting appropriate strategies before generating solutions.
Yet, this paradigm omits an implicit phase present in human cognition, where metacognitive monitoring precedes action. Before attempting complex tasks, humans – often without conscious deliberation – assess difficulty, retrieve relevant strategies, and establish confidence criteria (Flavell, 1979; Nelson and Narens, 1990). This metacognitive capacity operates by transforming world-centred uncertainty into self-centred propositional confidence, enabling both prospective planning and retrospective evaluation through mechanisms including global broadcast and post-decisional evidence accumulation (Fleming, 2024). Current reasoning architectures – whether Self-Verification (Weng et al., 2023) validating through backward verification of conclusions, SELF-REFINE (Madaan et al., 2023) improving through iterative feedback and refinement, or ReVISE (Lee et al., 2025) rethinking trajectories through intrinsic self-verification – lack these pre-generation assessment mechanisms, blindly committing to initial approaches that determine ultimate success or failure.
Although the Generate-Verify paradigm successfully mirrors Generation-Recognition models of memory recall (Bahrick, 1970; Kintsch, 1978), capturing how retrieval involves generation followed by recognition-based evaluation, these models specify only the mechanics of recall (object-level mechanics) – lacking the metacognitive control and monitoring that determines when and how these mechanics should be deployed (meta-level).
Cognitive psychology shows that successful task completion requires precisely such meta-level governance through monitoring task difficulty, selecting strategies, and deciding when to persist or abandon approaches (Flavell, 1979; Nelson and Narens, 1990). Flavell (1979) demonstrated that cognitive regulation emerges from dynamic interactions between metacognitive knowledge (beliefs about cognitive capabilities, task demands, and strategy effectiveness), metacognitive experiences (feelings of difficulty and comprehension), goals and strategies, with monitoring necessarily preceding action to enable appropriate strategy selection. Nelson and Narens (1990) further specified the hierarchical architecture of metacognition, distinguishing object-level processes from meta-level oversight that operates through control and monitoring flows. Their framework reveals how metacognitive experiences (Feeling of Knowing, Judgements of Learning) guide immediate decisions, with confidence thresholds dynamically adjusting through satisficing to balance accuracy against search costs.
Informed by these cognitive science foundations, we develop algorithmic formalisations that translate psychological theories into computational specifications suitable for reasoning systems. We formalise Flavell’s cognitive monitoring model and Nelson and Narens’ metamemory frameworks, translating them into computational algorithms to develop Monitor-Generate-Verify (MGV), a framework where explicit metacognitive monitoring precedes generation. Our formalisations translate abstract psychological constructs into algorithmic structures, highlighting gaps between human cognitive capabilities and current reasoning architectures.
@@
The remainder of this paper is organised as follows: Section 2 situates our contribution within current reasoning architectures and partial metacognitive implementations. Section 3 presents algorithmic formalisations of Flavell’s and Nelson and Narens’ theories, translating psychological concepts into computational structures – to our knowledge, the first systematic translation of these foundational theories into computational specifications. Section 4 demonstrates how these formalisations explain specific failures and suggest architectural improvements. Section 5 acknowledges implementation challenges, before Section 6 synthesises contributions and future directions.
@@ 2 Related Work
2.1 Metacognitive Capabilities in Language Models
Recent research explores metacognitive capabilities in LLMs through behavioural and neural analysis. Behavioural studies examine how models express uncertainty through confidence verbalisation (Wang et al., 2025; Tian et al., 2023; Xiong et al., 2023; Cash et al., 2024; Griot et al., 2025). Yet implicit confidence measures derived from token likelihoods demonstrate greater metacognitive sensitivity than explicitly prompted confidence statements (Xiong et al., 2023). This discrepancy between internal representation and external expression indicates that models access metacognitive signals they cannot adequately articulate – Lindsey et al. (2025) demonstrate that while Claude-3.5-Haiku accurately reports intermediate computational steps, it fabricates non-existent processes when queried about simple addition, despite correctly activating relevant neural mechanisms.
Direct neural analysis yields more substantive findings. Ji-An et al. (2025) identify a “metacognitive space” of lower dimensionality than models’ full neural space, indicating that LLMs monitor compressed representations rather than complete computational states. This observation aligns with established cognitive findings that metacognitive monitoring operates on abstracted representations (Reder, 1987). Furthermore, metacognitive signals such as confidence correspond to linearly separable directions in representation space (Zou et al., 2023; Liu et al., 2023), suggesting structured organisation of these internal monitoring capabilities.
2.2 Theoretical Foundations Gap
Despite the behavioural and neural evidence demonstrating LLM access to metacognitive signals – confidence representations, capability awareness, compressed monitoring spaces – such capabilities remain architecturally isolated. The distinction between signal detection and systematic control manifests in reasoning failures. Models experiencing the prefix dominance trap fail to revise suboptimal strategies (Luo et al., 2025), indicating that any metacognitive signals about task difficulty or strategy inadequacy remain functionally isolated from generation control.
The absence of systematic control architectures reflects a broader theoretical gap. Contemporary implementations incorporate metacognitive components as discrete modules rather than integrated systems. This fragmentation precludes the coordinated monitoring-control loops that Flavell identifies as necessary for strategy selection, and the hierarchical meta-level governance that Nelson and Narens demonstrate underlies successful cognitive regulation.
However, these psychological specifications lack computational translations. No principled framework exists for implementing such metacognitive architectures in artificial systems. This theoretical gap constrains both our understanding of reasoning failures and our capacity to develop systematic solutions. Without formal mappings from cognitive principles to algorithmic structures, implementations remain ad hoc, adopting surface features of metacognition while omitting the functional relationships that enable effective cognitive control.
Our contribution addresses this gap by providing algorithmic formalisations of established metacognitive theories. These formalisations preserve the functional dependencies identified in cognitive science while specifying implementation requirements for computational systems, thereby establishing theoretical foundations for metacognitive reasoning architectures.
@@ 3 Monitor-Generate-Verify (MGV)
Flavell (1979) and Nelson and Narens (1990) developed seminal theories of how metacognition coordinates cognitive processes through monitoring and control loops. These frameworks, though developed for human cognition, offer potential blueprints for computational systems. Flavell’s model provides a dynamic architecture where metacognitive knowledge and experience guide strategy selection and verification, while Nelson and Narens’ metamemory framework specifies how confidence thresholds and adaptive search mechanisms emerge from hierarchical monitoring and control. By computationally formalising these psychological theories, we establish Monitor-Generate-Verify (MGV) as a theoretical framework for understanding how explicit metacognitive mechanisms could address the architectural limitations of current reasoning systems. The following subsections present detailed formalisations that translate these cognitive science insights into algorithmic structures, revealing both what current architectures lack and how metacognitive principles might be operationalised computationally.
@@ 3.1 Flavell’s Model of Metacognition
Flavell (1979) conceptualises metacognition as a dynamic control architecture comprising four interacting components: metacognitive knowledge, metacognitive experience, goals (or tasks), and actions (or strategies). Rather than operating as independent modules, these components form an integrated system characterised by continuous bidirectional influences, positioning metacognition as a self-regulating system capable of adaptive control over cognitive processes. We present the core computational structure below, with a complete mathematical formalisation provided in Appendix A.
@@ 3.1.1 Cognitive Monitoring
Algorithm 1 Flavell’s Metacognitive Regulation
@@
The monitoring phase also generates metacognitive experiences of difficulty ($\mathcal{ME}_{\text{difficulty}}^{\tau}$), which Flavell (1979, p. 909) describes as subjective feelings of complexity, comprehension challenges, or sensing that material exceeds current capabilities. These experiences evolve through iterative assessments, progressing from initial coarse feelings to increasingly nuanced evaluations of specific challenge sources.
@@ 3.2 Nelson and Narens’ Model of Metamemory
Nelson and Narens (1990) establish metacognition as fundamentally hierarchical, distinguishing between object-level processes that operate on mental content and meta-level processes that operate on cognitive processes themselves. The meta-level maintains a dynamic internal representation of the object-level, enabling self-regulation through two distinct information flows: control (meta-level $\rightarrow$ object-level) and monitoring (object-level $\rightarrow$ meta-level). These relationships are logically independent and asymmetric – the meta-level maintains a model of the object-level while the object-level operates without corresponding meta-level representation. We present the core computational structure below, with complete mathematical formalisation provided in Appendix B.
@@ 3.2.1 Acquisition Process
Algorithm 2 Nelson and Narens: Acquisition
@@
The monitoring phase generates metacognitive experiences as multidimensional vectors. Ease of Learning (EOL) provides initial difficulty assessment, while Feeling of Knowing (FOK) incorporates prior outcomes to refine mastery judgements. These phenomenological experiences serve as primary input for control decisions (Nelson and Narens, 1990, p. 160). During the generation phase, resource allocation operates inversely to EOL/FOK values – items with lower metacognitive confidence receive proportionally more resources $r_{\tau,j}=R_{\text{total}}\times w_{j}/\sum_{k}w_{k}$ where $w_{j}=(\mathcal{ME}_{\tau,j}[1])^{-1}$. Strategy selection integrates these metacognitive inputs to map appropriate learning methods to individual items.
The verification phase employs Judgements of Learning (JOL) to evaluate mastery following cognitive outcomes. Items achieving the norm of study ($\text{JOL}_{\tau,j}\geq\mathcal{N}_{s}$) are removed from further consideration, whilst those below threshold remain in $\mathcal{J}_{\tau+1}$ for continued learning. The complete experience tuple accumulates in working memory as $\Phi_{\tau}^{\text{STM}}$, subsequently undergoing consolidation to LTM at encoding rate $\psi$.
@@ 3.2.2 Retrieval Process
Algorithm 3 Nelson and Narens: Retrieval

### v3

@@ Abstract
Test-time reasoning architectures such as those following the Generate-Verify paradigm, where a model iteratively refines or verifies its own generated outputs, prioritise generation and verification but exclude the monitoring processes that determine when and how reasoning should begin. This omission may contribute to the prefix dominance trap, in which models commit early to suboptimal reasoning paths and seldom recover, yielding roughly 20% accuracy loss. We address this architectural gap by proposing the Monitor-Generate-Verify (MGV) framework, a computational translation of Flavell’s and Nelson and Narens’ metacognitive theories that preserves their psychological detail. MGV extends the Generate-Verify paradigm by adding explicit monitoring that captures metacognitive experiences (from difficulty assessments to confidence judgements) before generation begins and refines future monitoring through verification feedback. Though we present no empirical validation, MGV provides a vocabulary for diagnosing component-level failures in reasoning systems, suggests specific architectural interventions for future designs, and identifies connections to resource-rational analysis that may ground its mechanisms in normative principles.
@@ 1 Introduction
Once language models commit to an initial reasoning strategy, subsequent verification rarely helps; this prefix dominance trap causes nearly 20% performance degradation when models choose suboptimal approaches, with limited recovery possible through refinement (Luo et al., 2025). Today’s dominant Generate-Verify test-time reasoning architectures (Weng et al., 2023; Madaan et al., 2023; Lee et al., 2025; Zhang et al., 2024) exemplify this limitation through their very design. They operate through immediate generation followed by iterative refinement, without assessing task characteristics or selecting appropriate strategies before generating solutions. What architectural vocabulary might help diagnose what these systems lack?
Cognitive psychology offers a candidate answer. Before attempting complex tasks, humans engage in metacognitive monitoring, assessing difficulty, retrieving relevant strategies, and establishing confidence criteria, often without conscious deliberation (Flavell, 1979; Nelson and Narens, 1990). This monitoring capacity transforms uncertainty into tractable signals that guide subsequent action. Flavell’s model specifies how metacognitive knowledge and experience interact to enable strategy selection, while Nelson and Narens’ metamemory framework details how confidence thresholds and adaptive search mechanisms emerge from hierarchical monitoring and control. These theories suggest that the Generate-Verify paradigm may be missing an entire phase of processing that precedes generation.
However, translating these psychological insights into architectural recommendations requires computational specification. Resource-rational analysis has made significant progress here, with Callaway et al. (2024) formalising Nelson and Narens’ metamemory theory as a meta-level MDP that derives optimal stopping policies from cost-benefit principles. While their approach offers normative grounding and empirical validation, the present work pursues a complementary goal: instead of deriving optimal behaviour from computational constraints, we translate the detailed psychological vocabulary of Flavell (1979) and Nelson and Narens (1990) into algorithmic form, preserving structure that resource-rational approaches may abstract away.
The resulting Monitor-Generate-Verify (MGV) framework extends the Generate-Verify paradigm by adding explicit monitoring that captures metacognitive experiences before generation begins and refines future monitoring through verification feedback. Our formalisation preserves constructs including the tripartite organisation of metacognitive knowledge, the distinction between difficulty and evaluative experiences, dual-counter evidence accumulation for feeling-of-knowing, and satisficing threshold dynamics. We offer no empirical validation and no normative justification for these architectural choices. The contribution is a detailed computational translation that provides diagnostic vocabulary for understanding what reasoning systems might lack, suggesting where architectural interventions might prove productive.
@@
The remainder of this paper is organised as follows. Section 2 situates our contribution within resource-rational approaches to metacognition and LLMs. Section 3 presents algorithmic formalisations of Flavell’s and Nelson and Narens’ theories, translating psychological concepts into computational structures. Section 4 acknowledges limitations and outlines future directions.
@@ 2 Related Work
This section situates the present work within two related research programmes. The first is resource-rational analysis, a cognitive modelling paradigm that formalises metacognition as cost-benefit optimisation over computational resources. The second is recent work applying these principles to language models. We review each in turn before clarifying how the present work differs in its foundational commitments.
2.1 Resource-Rational Analysis of Metacognition
Resource-rational analysis is a cognitive modelling paradigm that reconceives bounded rationality not as a departure from optimality but as optimality under constraints (Griffiths et al., 2015, 2019; Lieder and Griffiths, 2020). Making good decisions requires computation, yet computation itself is costly. This tension between the value of deliberation and its cost gives rise to a class of meta-problems: problems not about what to decide, but about how to decide. How much should an agent think before acting? Which cognitive strategy should be deployed for a given problem? Which aspects of a decision path merit exploration, and which can be safely ignored? These questions concern the allocation of computational resources, and answering them requires reasoning about reasoning itself.
Resource-rational analysis provides a framework for addressing such meta-problems by treating cognitive constraints as part of the optimisation problem rather than as departures from rationality. Under metalevel rationality, an agent is assessed not by the expected utility of actions taken, but by how well its cognitive algorithm trades off decision quality against deliberation costs. From this perspective, rationality concerns not only making good decisions but also employing efficient cognitive strategies. The framework reduces meta-reasoning [^1] Meta-reasoning should be distinguished from meta-learning. Meta-reasoning concerns how to allocate limited computational resources during cognition, deciding which computations to perform and when to stop deliberating. Meta-learning concerns how to structure the learning process itself, exploiting knowledge about the statistical structure of learning environments to make better use of limited data. Both address resource limitations, but at different levels: meta-reasoning optimises the use of time and cognitive effort within a task, whereas meta-learning optimises the acquisition of knowledge across tasks. For a treatment that situates both perspectives within the resource-rational framework, see Griffiths et al. (2019). to a decision-theoretic problem: for any computation the agent could perform, there exists a value of computation (VOC) that quantifies the expected improvement in decision quality minus the cost of deliberation.
This cognitive modelling paradigm provides a common structure that different research programmes have instantiated in distinct ways, each addressing a different facet of the meta-reasoning problem. Lieder and Griffiths (2017) modelled strategy selection: rather than asking whether to think more, this formulation asks how to think, treating the choice among cognitive strategies as a contextual bandit in which the agent learns to predict which mode of reasoning best suits each type of problem (Appendix 7.2). Lieder et al. (2018) modelled cognitive control: by casting control specification as a belief-MDP and introducing a learned approximation, this formulation explains how control exhibits plasticity, adapting through experience to select signals that optimally trade off effort against goal achievement (Appendix 7.4). Callaway et al. (2022) modelled planning: by embedding deliberation within a meta-level MDP, this formulation captures the sequential structure of mental simulation, recognising that the value of a computation depends not only on its immediate effect but also on the future computations it enables (Appendix 7.3). Callaway et al. (2024) modelled memory recall: by treating retrieval as evidence accumulation supervised by an optimal stopping process, this formulation reveals the adaptive function of feeling-of-knowing, providing the monitoring signal that enables efficient termination of searches unlikely to succeed while sustaining effort when retrieval remains probable (Appendix 7.5).
These four instantiations span a range of formal structures, from contextual bandits to belief-MDPs, and a range of cognitive phenomena, from high-level strategy choice to low-level memory retrieval. Yet they share a common theoretical commitment: the mind’s computational processes are themselves subject to rational optimisation. Appendix 7 provides a detailed technical treatment of these frameworks, comparing their formal structures and identifying shared algorithmic motifs. Here, we turn to recent work that has begun extending these principles beyond cognitive modelling to the training of language models.
2.2 Rational Metareasoning for Language Models
While the research reviewed above models human cognition, recent work has begun applying the VOC framework directly to language model training. De Sabbata et al. (2024) introduced RaM (Rational Metareasoning), which trains LLMs to use chain-of-thought reasoning selectively based on task difficulty. Their approach defines the reward of a reasoning chain $z$ as the difference between its utility and cost:
$R_{\pi}(x,y,z)=U_{\pi}(z|x,y)-C(z),$
where utility $U_{\pi}(z|x,y)=\log\pi_{\theta}(y|z,x)-\log\pi_{\theta}(y|x)$ quantifies the increase in likelihood of generating the correct answer $y$ when the chain of thought $z$ is included, and cost $C(z)=\gamma\cdot l(z)$ is proportional to the number of tokens. This formulation mirrors the equation, with reasoning tokens corresponding to intermediate computations and model outputs corresponding to actions. Models trained with this objective learn to generate shorter reasoning chains for easier problems and longer chains for harder problems, demonstrating that VOC-based training can induce adaptive computation in neural systems. This work provides empirical evidence that resource-rational principles can be operationalised in modern architectures, though it focuses on optimising existing reasoning capabilities rather than introducing new metacognitive components.
2.3 Relationship to the Present Work
Formalising metacognitive theory requires choosing what to treat as primitive and what to derive. Resource-rational analysis treats computational operations as primitive and derives metacognitive behaviour from cost-benefit optimisation. The present work makes the opposite choice, treating psychological constructs as primitive and translating them directly into algorithmic form. Put simply, instead of deriving that people think the way they do because it is efficient given limited cognitive resources, this work treats psychological ideas as the foundation and directly formalises them in computational terms.
This difference in approach leads to different contributions. Resource-rational analysis provides normative grounding, explaining why certain metacognitive policies are optimal given computational constraints. The present work provides psychological granularity, preserving detailed structure from the source theories that resource-rational models abstract away. Section 3 makes this concrete by formalising constructs including tripartite metacognitive knowledge, dual-counter evidence accumulation, and satisficing threshold dynamics. Section 4 returns to the relationship between these approaches and examines how they might complement one another.
@@ 3 Monitor-Generate-Verify (MGV)
Flavell (1979) and Nelson and Narens (1990) developed seminal theories of how metacognition coordinates cognitive processes through monitoring and control loops. These frameworks, though developed for human cognition, offer potential blueprints for computational systems. Flavell’s model provides a dynamic architecture where metacognitive knowledge and experience guide strategy selection and verification, while Nelson and Narens’ metamemory framework specifies how confidence thresholds and adaptive search mechanisms emerge from hierarchical monitoring and control. By computationally formalising these psychological theories, we establish Monitor-Generate-Verify (MGV) as a theoretical framework for understanding how explicit metacognitive mechanisms could address the architectural limitations of current reasoning systems. The following subsections present detailed formalisations that translate these cognitive psychology insights into algorithmic structures, revealing both what current architectures lack and how metacognitive principles might be operationalised computationally.
@@ 3.1 Flavell’s Model of Metacognition
Flavell (1979) conceptualises metacognition as a dynamic control architecture comprising four interacting components: metacognitive knowledge, metacognitive experience, goals (or tasks), and actions (or strategies). Rather than operating as independent modules, these components form an integrated system characterised by continuous bidirectional influences, positioning metacognition as a self-regulating system capable of adaptive control over cognitive processes. We present the core computational structure below, with a complete mathematical formalisation provided in Appendix 5.
@@ 3.1.1 Cognitive Monitoring
Algorithm 1 Flavell’s Model of Cognitive Control
@@
The monitoring phase also generates metacognitive experiences of difficulty ($\mathcal{ME}_{\text{difficulty}}^{\tau}$), which Flavell (1979, p. 909) describes as subjective feeling-of-complexity, comprehension challenges, or sensing that material exceeds current capabilities. These experiences evolve through iterative assessments, progressing from initial coarse feelings to increasingly nuanced evaluations of specific challenge sources.
@@ 3.2 Nelson and Narens’ Model of Metamemory
Nelson and Narens (1990) establish metacognition as fundamentally hierarchical, distinguishing between object-level and meta-level processes. The object-level comprises the cognitive operations themselves, such as searching memory for a target item. The meta-level supervises these operations, by monitoring progress toward the goal (e.g., how close is recall?) and controlling how long the process is allowed to continue (e.g., should the search persist or terminate?). This supervision relies on two distinct information flows: monitoring conveys information upward from object-level to meta-level, while control conveys decisions downward from meta-level to object-level. These relationships are logically independent and asymmetric: the meta-level maintains a model of the object-level, while the object-level operates without corresponding meta-level representation. We present the core computational structure below, with complete mathematical formalisation provided in Appendix 6.
@@ 3.2.1 Acquisition Process
Algorithm 2 Nelson and Narens’ Model of Acquisition
@@
The monitoring phase generates metacognitive experiences as multidimensional vectors. Ease-of-learning (EOL) provides initial difficulty assessment, while feeling-of-knowing (FOK) incorporates prior outcomes to refine mastery judgements. These phenomenological experiences serve as primary input for control decisions (Nelson and Narens, 1990, p. 160). During the generation phase, resource allocation operates inversely to EOL/FOK values – items with lower metacognitive confidence receive proportionally more resources $r_{\tau,j}=R_{\text{total}}\times w_{j}/\sum_{k}w_{k}$ where $w_{j}=(\mathcal{ME}_{\tau,j}[1])^{-1}$. Strategy selection integrates these metacognitive inputs to map appropriate learning methods to individual items.
The verification phase employs judgement-of-learning (JOL) to evaluate mastery following cognitive outcomes. Items achieving the norm of study ($\text{JOL}_{\tau,j}\geq\mathcal{N}_{s}$) are removed from further consideration, while those below threshold remain in $\mathcal{J}_{\tau+1}$ for continued learning. The complete experience tuple accumulates in working memory as $\Phi_{\tau}^{\text{STM}}$, subsequently undergoing consolidation to LTM at encoding rate $\psi$.
@@ 3.2.2 Retrieval Process
Algorithm 3 Nelson and Narens Model of Retrieval